on the 50 foot side of a house, a 3 sided fence is made out of 44 feet of fencing, with 2 of the three sides being the same length.

What is the maximum area, in square feet for the garden if all of the 44 feet of fencing is used
Answer #1

fencing = 44 feet

let length = l

width = w

2w + l = 44

l = 44- 2w

area = length * width

= w( 44 - 2w )

= 44w - 2w^2

maximum area occurs at

w = - b/2a

w = - 44/ 2(-2)

w = 11

l = 44 - 2(11) = 22

area = 22(11)

maximum area = 242 square feet